. A town of 2,000 households constitutes a market for eggs. Current sales are 2400 dozen eggs per week at a price of $1.25 per dozen. 1200 households living on the west side of the river buy1600 dozen eggs and their elasticity of demand is -1.5. The remaining households live on the east side of the river, buy the rest of the eggs and have an elasticity of demand of -3. Calculate the elasticity of market demand curve for the town as a whole.  



1
Expert's answer
2021-04-19T07:41:32-0400

"800\times 1.25=1000"

"1200\times 1.25 = 1500"

"(1000−1500)\times [\frac{1000+15000}2]\times 100"

"=\frac{−500}{1250}\times 100"

"800\u22121200=\u2212400"

"800+1200 = 2000"

"\frac{-400}{2000}\times*100 = -20"

"\frac{-20}{40} = -0.5"

PED of market demand curve for the town as a whole = -0.5